# DiFSD: Ego-Centric Fully Sparse Paradigm with Uncertainty Denoising and Iterative Refinement for Efficient End-to-End Self-Driving

### v4

@@ 3.7 End-to-End Learning
where is a combination of binary classification loss and regression loss to learn geometric score, where the positive (interactive) samples are denoted as grid cells with geometric score (within 3 for each future waypoint). An additional regression loss is included in for ego status prediction, instead of directly using it as input to the planner as [11, 39, 17], which will lead to information leakage as proven in [21]. Meanwhile, vectorized planning constrains identified in [17] such as collision, overstepping and direction are also included in for regularization. is the number of motion planning stages.
@@ 4.2 Implementation Details
DiFSD plans a 3 future ego-trajectory with 2 history information as input. Our DiFSD has two variants, namely DiFSD-S and DiFSD-B. As for DiFSD-S, both perspective-view perception version DiFSD-S (PV) and BEV perception version DiFSD-S (BEV) are all implemented for comparison. ResNet50 [8] is adopted as the default backbone network for visual encoding. The perception range is set to 6030 longitudinally and laterally. Input image size of DiFSD-S is resized to . For DiFSD-S (BEV), the default number of BEV query, map query, agent query is , and 300, respectively. For DiFSD-S (PV), is 6, is 900 and is 100 respectively. Each map element contains 20 map points. The feature dimension is 256. The noise scale is set to 2.0 and 0.2FD for motion and planning respectively. is set to 3. DiFSD-B has larger input image resolution ( and backbone network (ResNet101). We use AdamW [30] optimizer and Cosine Annealing [29] scheduler to train DiFSD with weight decay 0.01 and initial learning rate -4. DiFSD is trained for 60 epochs and 4 epochs on nuScenes and Bench2Drive respectively, running on 8 NVIDIA Tesla A100 GPUs with batch size 8 empirically.
@@ 4.3 Main Results
Open-loop Planning Evaluation. As show in Tab. 1, DiFSD shows great advantages in both performance and efficiency compared with previous works. On one hand, DiFSD-S achieves the minimum L2 error even with lightweight visual backbone and inferior BEV perception encoder. Specifically, compared with BEVFormer-based end-to-end methods [11, 17], DiFSD-S (BEV) reduces the average L2 error by a great margin (0.38 and 0.37, separately), while significantly reducing the average collision rates by 84% and 52% respectively. Equipped with deeper visual backbone and advanced sparse detectors from Perspective View (PV), the average L2 error and collision rates can be further reduced to 0.32 and to 0.06 respectively. Notably, we are the first to achieve 0 collision rate on 1. On the other hand, benefiting from the ego-centric hierarchical interaction, only sparse interactive agents (2%) are considered for motion planning. Hence, DiFSD-S can achieve great efficiency with 14.8 FPS, and faster than UniAD [11] and VAD [17] respectively.
Closed-loop Planning Evaluation. We further validate the closed-loop performance in Bench2Drive [15], which has been proposed recently for comprehensive benchmarking of end-to-end planning methods. As shown in Tab. 2, AD-MLP [41] has a high L2 error and bad closed-loop planning performance using merely ego status as input, which is different from findings in nuScenes [1], demonstrating the behavior diversity in Bench2Drive. UniAD [11] has a lower L2 error compared to VAD [17] but with worse closed-loop planning performance as discussed in [21]. Notably, DiFSD-S achieves both the lowest L2 error and best closed-loop performance with great efficiency, showcasing the superiority and generalizability of our proposed method.
@@ 4.4 Ablation Study
We conduct extensive experiments to study the effectiveness of our DiFSD. We use DiFSD-S as default for ablation. More experiments are provided in the Appendix.
Necessity of Geometric Prior. We claim that the Closest In-Path Vehicle as well as Stationary (CIPV / CIPS) are more likely to interact with the ego-vehicle. To verify the necessity of such geometric prior, we conduct exhaustive ablations of the ego-centric query selector as show in Tab. 3. Without ego-centric selection, fewer objects randomly selected can result in worse planning results. While using the ego-centric cross attention, only 2% of surrounding queries are enough for achieving convincing planning performance. Besides, introducing the geometric prior through attention can dramatically reduce the L2 error and collision rate by 8% and 42% respectively. Meanwhile, when utilizing the ground-truth geometric score for upper-limit evaluation, we can obtain the extremely excellent planning performance (0.23 average L2 error). Undoubtedly, the proposed ego-centric query selector equipped with geometric attention is nontrivial for motion planner.
Effect of designs in Hierarchical Interaction. Tab. 4 shows the effectiveness of our elaborate designs in the hierarchical interaction module, which contains three main designs such as Dual Interaction (DI), Geometric Attention (GA) and Coarse-to-Fine Selection (CFS). DI models both ego-centric and object-centric interactions respectively, which improves the planning performance greatly as expected. GA facilitates the query selection process as discussed in Tab. 3, which reduces the collision rate by a great margin (42%). And CFS contributes to the interaction modeling quality through hierarchical receptive fields from global to local. All of these three designs combined together can achieve overall convincing planning performance.
@@
Effect of designs in Motion Planner. As for motion planner in DiFSD, Joint Motion Prediction (JMP), Planning Optimization (PO) as well as Iterative Refinement (IR) makes up the planning pipeline of ego-vehicle. Besides, Uncertain Denoising (UD) contributes to the system stability and training convergence. Tab. 5 explores the effect of each design exhaustively. ID-1 indicates evaluating the proposal trajectory of ego-vehicle predicted together with interactive agents, which achieves competitive L2 error but is easier to collide with surrounding agents. ID-2 improves the collision rate greatly by 38.9% with the help of PO and planning constraints [17] during training phase. ID-4 emphasizes the importance of IR in improving the quality of ego-planning trajectory (average 5.4% L2 error and 36.3% collision rate reduction respectively). ID-3 reflects the benefit of UD used for end-to-end training compared to ID-4.
Effect of Iterative Refinement stages. We continue to study the number of refinement stages in Tab. 6. We can observe that our DiFSD can obtain superior planning performance with one additional refinement stage (36.3% collision rate reduction), which becomes saturated when introducing more stages. Hence, two-stage interacted motion planner is enough for achieving convincing results.
@@
Runtime of each module. We evaluate the modular runtime of DiFSD-S with different perception paradigms as shown in Tab. 8. Generally, visual backbone and perception encoder occupy the most of the runtime for feature extraction and scene understanding. Hierarchical interaction also takes a significant part for interaction modeling and interactive query selection. Thanks to the ego-centric interaction and selection module, the joint motion planner only consumes 7.9ms to plan the future ego-trajectory. Notably, though PV-based perception module [24] achieves superior detection performance, it consumes more computation cost due to the repeated query projection, deformable feature aggregation as well as the symmetric sparse perception architecture for unified object detection and online mapping, without a shared BEV feature as [17, 11].
4.5 Qualitative Results
We visualize the motion trajectories of sparse interactive agents as well as planning results of DiFSD as illustrated in Fig. 5. Both surrounding camera images and prediction results on BEV are provided accordingly. Besides, we also project the planning trajectories to the front-view camera image. Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of DiFSD motivation. DiFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.
@@ 4.5 Qualitative Results
| Module | Bird’s-Eye View (BEV) Perception | Perspective-View (PV) Perception | ||
| Latency (ms) | Proportion (%) | Latency (ms) | Proportion (%) | |
| Backbone | 8.4 | 12.4 | 11.0 | 11.8 |
| Perception Encoder | 34.4 | 50.8 | 54.9 | 58.6 |
| Hierarchical Interaction | 17.0 | 25.1 | 19.9 | 21.2 |
| Motion Prediction | 4.5 | 6.6 | 4.5 | 4.8 |
| Planning Optimization | 3.4 | 5.1 | 3.4 | 3.6 |
| Total | 67.7 | 100 | 93.7 | 100 |
@@ 5 Conclusion
In this paper, we propose a fully sparse paradigm for end-to-end self-driving in an ego-centric manner, termed as DiFSD. DiFSD revisits the human driving behavior and conducts hierarchical interaction based on sparse representation and perception results. Only interactive agents are considered for joint motion prediction with the ego-vehicle. Iterative planning optimization strategy contributes to the driving safety with high efficiency. Besides, uncertainty modeling is conducted to improve the stability of end-to-end system. Extensive ablations and comparisons reveal the superiority and great potential of our ego-centric fully sparse paradigm for future research.
@@ Appendix A Evaluation Metrics
Perception. The evaluation for detection and tracking follows standard evaluation protocols [1]. For detection, we use mean Average Precision(mAP), mean Average Error of Translation(mATE), Scale(mASE), Orientation(mAOE), Velocity(mAVE), Attribute(mAAE) and nuScenes Detection Score(NDS) to evaluate the model performance. For online mapping, we calculate the Average Precision(AP) of three map classes: lane divider, pedestrian crossing and road boundary, then average across all classes to get mean Average Precision (mAP).
Planning. We adopt commonly used L2 error and collision rate to evaluate the planning performance. The evaluation of L2 error is aligned with VAD [17]. For collision rate, there are two drawbacks in previous [11, 17] implementation, resulting in inaccurate evaluation in planning performance. On one hand, previous benchmark convert obstacle bounding boxes into occupancy map with a grid size of 0.5m, resulting in false collisions in certain cases, e.g. ego vehicle approaches obstacles that smaller than a single occupancy map pixel [41]. (2) The heading of ego vehicle is not considered and assumed to remain unchanged [21]. To accurately evaluate the planning performance, we account for the changes in ego heading by estimating the yaw angle through trajectory points, and assess the presence of a collision by examining the overlap between the bounding boxes of ego vehicle and obstacles. We reproduce the planning results on our benchmark with official checkpoints [11, 17] for a fair comparison.
Appendix B More Ablation Study
@@
| DiFSD-S | ResNet50 | ✓ | 32.8 | 45.8 |
| DiFSD-S | ResNet50 | ✗ | 41.0 | 52.8 |
| DiFSD-B | ResNet101* | ✗ | 49.6 | 58.9 |
@@
| DiFSD-S (BEV) | 46.7 | 54.3 | 56.0 | 52.3 |
| DiFSD-S (PV) | 54.9 | 55.7 | 57.3 | 56.0 |
| DiFSD-B (PV) | 52.3 | 58.2 | 59.3 | 56.6 |
Appendix C Analysis & Discussion
@@
The GroundTruth future state distribution of ego-vehicle on nuScenes validation set is illustrated in Fig. A1, which is calculated with fixed time interval (1) between consecutive predicted waypoints. And we also compare the output ego-state distribution of different popular end-to-end methods based on planned trajectories respectively, as show in Fig. A2. We can observe that without ego-centric design, the optimized end-to-end model is unable to handle various emergencies appearing in the driving scenarios, where the absolute values of and are larger than normal situations. Under this circumstance, the output planned trajectories cannot conform to the expert routes as expected. However, our DiFSD performs consistently better in planning the future ego states with variable speed and acceleration, owing to the ego-centric hierarchical interaction and selection mechanism, thus the iterative motion planner can focus on the interactive agents rather than irrelevant objects.
Appendix D Visualization
As show in Fig. A3-A5, we provide more visualization results to illustrate the effectiveness of DiFSD on various driving scenarios (i.e., interactive / non-interactive scenes, lane-change / lane-keep / following scenes, overtaking / avoidance scenes, intersection scenes) under different commands (i.e., “Go Straight”, “Turn Left”, “Turn Right”).

### v5

@@ 3.7 End-to-End Learning
where is a combination of binary classification loss and regression loss to learn geometric score, where the positive (interactive) samples are denoted as grid cells with geometric score (within 3 for each future waypoint). An additional regression loss is included in for ego status prediction, instead of directly using it as input to the planner as [12, 47, 17], which will lead to information leakage as proven in [23]. Meanwhile, vectorized planning constrains identified in [17] such as collision, overstepping and direction are also included in for regularization. Besides, the weight terms of different losses are adjusted empirically to ensure the same magnitude. is the number of motion planning stages.
@@ 4.2 Implementation Details
EgoFSD plans a 3 seconds future ego-trajectory with 2 seconds history information as input, which has two variants, namely EgoFSD-S and EgoFSD-B. As for EgoFSD-S, both Perspective-View version and BEV version are all implemented. ResNet50 [8] is adopted as the default backbone network for visual encoding. The perception range is set to 6030 longitudinally and laterally. Input image size of EgoFSD-S is resized to . For EgoFSD-S (BEV), is 3, and the number of BEV query, map query, agent query are , and 300, respectively. For EgoFSD-S (PV), is 900 and is 100 respectively. Each map element contains 20 map points. The feature dimension is 256. The noise scale is set to 2.0 and 0.2FD for motion and planning respectively. is set to 3. EgoFSD-B has deeper network (ResNet101, = 6) and larger image resolution (. We use AdamW [33] optimizer and Cosine Annealing [32] scheduler to train EgoFSD with weight decay 0.01 and initial learning rate -4. EgoFSD is trained for 60 epochs and 6 epochs on nuScenes and Bench2Drive respectively, running on 8 NVIDIA Tesla A800 GPUs with total batch size 32 empirically.
| Method | Open-loop Metric | Closed-loop Metrics | ||
| Avg. L2 (m) | Driving Score | Success Rate (%) | Efficiency | |
| AD-MLP [49] | 3.64 | 18.05 | 0.00 | 48.45 |
| UniAD-Tiny [12] | 0.80 | 40.73 | 13.18 | 123.92 |
| UniAD-Base [12] | 0.73 | 45.81 | 16.36 | 129.21 |
| VAD [17] | 0.91 | 42.35 | 15.00 | 157.94 |
| GenAD [52] | - | 44.81 | 15.90 | - |
| MomAD [38] | 0.82 | 44.54 | 16.71 | 170.21 |
| EgoFSD-S | 0.70 | 52.02 | 21.00 | 178.30 |
| EgoFSD-B | 0.66 | 60.39 | 31.78 | 180.63 |
@@ 4.3 Main Results
Open-Loop Planning Evaluation. As shown in Tab. 1, EgoFSD demonstrates significant advantages in both performance and efficiency compared to previous works. On one hand, EgoFSD-S achieves the minimum L2 error despite its lightweight visual backbone and inferior BEV perception. Specifically, compared with BEVFormer-based end-to-end methods [12, 17], EgoFSD-S (BEV) reduces the average L2 error by a great margin (0.38 and 0.37, separately), while significantly reducing the average collision rates by 84% and 52% respectively. Equipped with deeper visual backbone and advanced sparse detectors from Perspective View (PV), the average L2 error and collision rates can be further reduced to 0.30 and to 0.06 respectively. On the other hand, benefiting from the ego-centric hierarchical interaction, only sparse interactive agents (2%) are considered for motion planning. Hence, EgoFSD-S can achieve great efficiency with 16.7 FPS, and faster than UniAD [12] and VAD [17] respectively.
Closed-Loop Planning Evaluation. We further validate the closed-loop performance in Bench2Drive [15], which has been proposed recently for comprehensive benchmarking of end-to-end planning methods. As shown in Tab. 2, AD-MLP [49] has a high L2 error and bad closed-loop planning performance using merely ego status as input, which is different from findings in nuScenes [1], demonstrating the behavior diversity in Bench2Drive. UniAD [12] has a lower L2 error compared to VAD [17] but with worse closed-loop planning performance as discussed in [23]. Notably, EgoFSD-S achieves both the lowest L2 error and best closed-loop performance with great efficiency, showcasing the superiority and generalizability of our proposed method.
@@ 4.4 Ablation Study
We conduct extensive experiments to study the effectiveness of our EgoFSD. We use EgoFSD-S as default for ablation. More experiments are provided in the Appendix.
Necessity of Geometric Prior. We claim that the Closest In-Path Vehicle as well as Stationary (CIPV / CIPS) are more likely to interact with the ego-vehicle. To verify the necessity of such geometric prior, we conduct exhaustive ablations of the ego-centric query selector as show in Tab. 3. Without ego-centric selection, fewer objects randomly selected can result in worse planning results. While using the ego-centric cross attention, only 2% of surrounding queries are enough for achieving convincing planning performance. Besides, introducing the geometric prior through attention can dramatically reduce the collision rate by 42% especially, thanks to the superior interactive planning. Meanwhile, when utilizing the Ground-Truth geometric score for upper-limit evaluation, we can obtain the extremely excellent planning performance (0.23 average L2 error). The proposed ego-centric query selector equipped with geometric attention is nontrivial for motion planner.
Effect of designs in Hierarchical Interaction. Tab. 4 shows the effectiveness of our elaborate designs in the hierarchical interaction module, which contains three main designs such as Dual Interaction (DI), Geometric Attention (GA) and Coarse-to-Fine Selection (CFS). DI models both ego-centric and object-centric interactions respectively, which ensures the query selection quality for interactive planning. GA facilitates the query selection process as discussed in Tab. 3, which reduces the collision rate obviously owing to the specialized attention on CIPV / CIPS. And CFS contributes to the multi-granularity interaction modeling through hierarchical receptive fields from global to local. All of these three designs combined together can achieve overall convincing planning performance.
@@
Effect of designs in Motion Planner. As for motion planner in EgoFSD, Joint Motion Prediction (JMP), Planning Optimization (PO) as well as Iterative Refinement (IR) makes up the planning pipeline of ego-vehicle. Besides, Uncertain Denoising (UD) contributes to the system stability and training convergence. Tab. 5 explores the effect of each design exhaustively. ID-1 indicates evaluating the proposal trajectory of ego-vehicle predicted together with interactive agents, which achieves competitive L2 error but is easier to collide with surrounding agents. ID-2 improves the collision rate greatly by 38.9% with the help of PO and planning constraints [17] during training phase. ID-4 emphasizes the importance of IR in improving the quality of ego-planning trajectory (average 5.4% L2 error and 36.3% collision rate reduction respectively). ID-3 reflects the benefit of UD used for end-to-end training compared to ID-4.
Effect of Iterative Refinement stages. We continue to study the number of refinement stages in Tab. 6. We can observe that our EgoFSD can obtain superior planning performance with one additional refinement stage (36.3% collision rate reduction), which becomes saturated when introducing more stages. Hence, two-stage interacted motion planner is enough for achieving convincing results.
@@
Runtime of each module. We evaluate the modular runtime of EgoFSD-S with Perspective-View perception paradigms as shown in Tab. 8. Generally, visual backbone and perception decoder occupy the most of the runtime (57.7%) for feature extraction and scene understanding. Hierarchical interaction also takes a significant part for interactive query selection. Thanks to the ego-centric interaction and selection module, the joint motion planner only consumes 7.5ms to plan the future ego-trajectory.
4.5 Qualitative Results
We visualize the motion trajectories of sparse interactive agents as well as planning results of EgoFSD as illustrated in Fig. A3. Both surrounding camera images and prediction results on BEV are provided accordingly. Besides, we also project the planning trajectories to the front-view camera image. Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of EgoFSD motivation. EgoFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.
| Module | #Param (M) | Latency (ms) | Proportion (%) |
| Backbone | 26.7 | 4.7 | 7.9 |
| Sparse Perception | 11.2 | 29.8 | 49.8 |
| Hierarchical Interaction | 21.3 | 17.8 | 29.8 |
| Joint Motion Prediction | 4.9 | 4.1 | 6.9 |
| Planning Optimization | 1.8 | 3.4 | 5.6 |
| Total | 65.9 | 59.8 | 100.0 |
@@ 5 Conclusion
In this paper, we propose a Fully Sparse Paradigm for end-to-end self-driving in an Ego-Centric manner, termed as EgoFSD. EgoFSD conducts hierarchical interaction based on sparse representation and perception results. Only interactive agents are considered for joint motion prediction including the ego-vehicle. Iterative planning optimization strategy contributes to the driving safety with interactive decision. Besides, uncertainty modeling is conducted to improve the stability of end-to-end system. Extensive ablations and comparisons reveal the superiority and great potential of our ego-centric fully sparse paradigm.
@@ Appendix A Evaluation Metrics
Perception. The evaluation for detection and online mapping follows standard evaluation protocols [1]. For detection, we use mean Average Precision (mAP), mean Average Error of Translation (mATE), Scale (mASE), Orientation (mAOE), Velocity (mAVE), Attribute (mAAE) and nuScenes Detection Score (NDS) to evaluate the model performance. For online mapping, we calculate the Average Precision (AP) of three map classes: lane divider, pedestrian crossing and road boundary, then average across all classes to get mean Average Precision (mAP).
Open-Loop Planning. We adopt commonly used L2 error and collision rate to evaluate the planning performance. The evaluation of L2 error is aligned with VAD [17]. For collision rate, there are two drawbacks in previous [12, 17] implementation, resulting in inaccurate evaluation in planning performance. On one hand, previous benchmark convert obstacle bounding boxes into occupancy map with a grid size of 0.5m, resulting in false collisions in certain cases, e.g. ego vehicle approaches obstacles that smaller than a single occupancy map pixel [49]. (2) The heading of ego vehicle is not considered and assumed to remain unchanged [23]. To accurately evaluate the planning performance, we account for the changes in ego heading by estimating the yaw angle through trajectory points, and assess the presence of a collision by examining the overlap between the bounding boxes of ego vehicle and obstacles. We reproduce the planning results on our benchmark with official checkpoints [12, 17] for a fair comparison.
Closed-Loop Planning. We use the official 220 routes and official metrics of Bench2Drive [15] for evaluation. The Driving Score (DS) is defined as the product of Route Completion and Infraction Score, capturing both task completion and rule adherence. The Success Rate (SR) measures the percentage of routes completed successfully within the allocated time. A route is deemed successful if the ego vehicle reaches its destination without any traffic rule violation. Efficiency quantifies the vehicle’s velocity relative to surrounding traffic, encouraging progressiveness without aggression.
Appendix B More Details
Problem Formulation. Given multi-view camera image sequence can be denoted as , where is the number of camera views and indicates the temporal length till current timestep respectively. Annotation of input for end-to-end planning is composed by a set of future waypoints of the ego-vehicle , where = 3 seconds (2Hz) is the planning time horizon, and (, ) is the BEV location transformed to the ego-vehicle coordinate system at current timestep . Meanwhile, driving command as well as ego-status is also provided. Annotation set is used during training. During prediction, the planned trajectory of ego-vehicle should fit the annotation with minimum L2 errors and collision rate with surrounding agents.
Data Augmentation. During training phase, we conduct a global data augmentation strategy to improve the model stability. Specifically, random rotation, translation and flipping (Y-axis) are applied to both bounding boxes of agent / map instances and future trajectories of agents / ego-vehicle respectively. Meanwhile, we adjust the camera extrinsic parameters accordingly to ensure spatial alignment between the surround-view images and the noise-perturbed targets.
Appendix C More Ablation Study
@@
| EgoFSD-S (BEV) | ResNet50 | ✓ | 32.8 | 45.8 |
| EgoFSD-S (PV) | ResNet50 | ✗ | 41.0 | 52.8 |
| EgoFSD-B (PV) | ResNet101* | ✗ | 49.6 | 58.9 |
@@
| EgoFSD-S (BEV) | 46.7 | 54.3 | 56.0 | 52.3 |
| EgoFSD-S (PV) | 54.9 | 55.7 | 57.3 | 56.0 |
| EgoFSD-B (PV) | 52.3 | 58.2 | 59.3 | 56.6 |
Appendix D Effectiveness Analysis
@@
The Ground-Truth future state distribution of ego-vehicle on nuScenes validation set is illustrated in Fig. A1, which is calculated with fixed time interval (1 second) between consecutive output future waypoints. And we also compare the output ego-state distribution of different popular E2E methods based on planned trajectories respectively, as show in Fig. A2. We can observe that without ego-centric design, existing end-to-end models are incapable of handling various emergencies appearing in the driving scenarios, where the absolute values of and are larger than normal situations (i.e., emergent braking, acceleration, turning or obstacle avoiance). Under this circumstance, the output planned trajectories cannot conform to the recorded expert demonstrations as expected. However, our EgoFSD performs consistently better in planning the future trajectory with dynamic and adaptive ego states of variable speed and acceleration. Owing to the ego-centric hierarchical interaction and selection mechanism, the iterative motion planner can focus on the interactive agents rather than general objects.
Appendix E More Discussion
Practicality and Generalizability of Ego-Centric Design. We empirically validate the necessity and effectiveness of our proposed ego-centric query selection and interactive planning with the help of geometric prior in Sec. 4.4. Moreover, we claim that the usage of geometric prior through interactive score fusion is general and practical for data-driven E2E-AD training, which considers both semantic relevance of driving context and human-like driving attention distribution according to current driving situation and future goal waypoint provided by the route planner. Extensive experiments conducted on both popular nuScenes [1] and large-scale challenging Bench2Drive [15] benchmark demonstrate the generalizability and significant improvement of computational efficiency.
Comparison to Diffusion-based Planners. We notice that recent works [26, 41] also explore diffusion model for end-to-end autonomous driving application. DiffusionDrive [26] learns denoising from anchored Gaussian distribution to the multi-mode driving action distribution with a truncated diffusion policy. DiffAD [41] treats E2E-AD as a conditional image generation task, which adopts diffusion model to predict noise from the noisy latent BEV image. Differently, we conduct a two-level data augmentation strategy to enrich the training samples including position diffusion of detected agents and trajectory diffusion of proposal ego trajectory. In this way, the joint motion prediction result can be more robust even with uncertain agent positions. And the offset prediction of ego trajectory can be more accurate in the planning optimization stage. Extensive ablations demonstrate the effectiveness of our uncertainty denoising design in improving the performance and training stability of EgoFSD.
Comparison to other Sparse Frameworks. Prior to our work, some exiting methods [39, 50] are also built upon sparse representation for end-to-end autonomous driving. SparseDrive [39] introduces the symmetric sparse perception architecture for object detection and online mapping respectively. However, it consumes more computation cost due to the repeated query projection, deformable feature aggregation, without a shared BEV feature as [17, 12]. Our EgoFSD constructs a shared 3D position-aware feature as [7] to facilitate parallel sparse query decoding for perception and planning. Besides, an additional core contribution of our EgoFSD lies on an ego-centric design, which aims to select the most important agent/map instances (i.e., CIPV/CIPS) for interactive modeling. Benefiting from the geometric prior and iterative refinement, the query selection quality can be gradually enhanced for efficient planning. Extensive ablations showcase the superiority of our proposed Ego-Centric Fully Sparse Paradigm.
Appendix F Limitations and Social Impact
Limitations. There still remains some limitations in our work. First, EgoFSD performs ego-centric query selection based on the geometric score, which is obtained through referring to the distance map using static query positions. And the motion information of surrounding agents are introduced into the ego-object dual interaction layer implicitly through iteratively updating the agent queries. How to generate better geometric score considering agent dynamics explicitly is worthy of further discussion. Besides, how to incorporate additional traffic signals and vision-language models into current driving system for autonomous decision making also deserves further exploration.
Social Impact. EgoFSD could be easily deployed on mass-produced car chips with different limits of computing resources, and thus can be served as a plug-and-play software to assist human drivers in decision-making and safe driving.
Appendix G Qualitative Results
We visualize the motion trajectories of sparse interactive agents as well as planning results of EgoFSD as illustrated in Fig. A3. Both surrounding camera images and prediction results on BEV are provided accordingly. Besides, we also project the planning trajectories to the front-view camera image. Only the top-3 trajectories of selected agents interacting with ego-vehicle are visualized for better understanding of EgoFSD motivation. EgoFSD outputs planning results based on the vectorized representation in an end-to-end manner, not requiring any dense interaction and redundant motion modeling, let alone hand-crafted post-processing.